# Super approximation for $\text{SL}_{2}\times\text{SL}_{2}$ and $\text{ASL}_{2}$

### v3

@@ 1. Introduction
For a set $\mathcal{A}$ of positive integers, we say $G$ has super approximation with respect to $\mathcal{A}$ if there is $\epsilon>0$ such that $\lambda_{0,q}-\lambda_{1,q}>\epsilon$, $\forall q\in\mathcal{A}$. The existence of a uniform positive spectral gap only depends on $G$ and is independent of the choice of the finite generating set $S$. If $\mathcal{A}=\mathbb{Z}_{+}$, we simply say $G$ has super approximation.
It is well known that a large class of lattices in $\text{GL}_{n}(\mathbb{Z})$ satisfy the super approximation property, a discovery originally due to Margulis [Mar73] and enhanced later by [BS91] and [Clo03], but the involved techniques (spectral method, Property $T$, automorphic forms, etc.) only works for arithmetic lattices and can not deal with a general non-lattice discrete group. A breakthrough came from Bourgain and Gamburd, who developed the so-called “Bourgain-Gamburd expansion machine”, which is an analytic-combinatorial tool that allows them to prove super approximation property for any Zariski-dense subgroup of $\text{SL}_{2}(\mathbb{Z})$ with respect to prime moduli [BG08b]. A critical ingredient in the proof is Helfgott’s triple product theorem [Hel08]. Since then, there has been a series of papers extending Bourgain-Gamburd’s Theorem to more general groups with respect to more general moduli [BG08a], [BG09], [BGS10], [GV12], [Var12], [BV12], [Gol19], [HdS21].
@@
Let $G<{\text{SL}}_{n}(\mathbb{Z})$ be finitely generated. Then $G$ has the super approximation property if and only if the identity component $\mathbb{G}_{0}$ of the Zariski closure $\mathbb{G}$ of $G$ is perfect, i.e. $[\mathbb{G}_{0},\mathbb{G}_{0}]=\mathbb{G}_{0}$.
@@
(Salehi-Golsefidy)[Gol19] Let $G<{\text{SL}}_{n}(\mathbb{Z})$ be finitely generated, then $G$ has the super approximation property with respect to bounded powers of square free integers coprime to $q_{0}$ if and only if the identity component $\mathbb{G}_{0}$ of the Zariski closure $\mathbb{G}$ of $G$ is perfect, i.e. $[\mathbb{G}_{0},\mathbb{G}_{0}]=\mathbb{G}_{0}$.
In the paper [Gol19], the author also claims super approximation for $G$ over $\{p^{n}\}_{n\in\mathbb{Z}_{+}}$ for a fixed prime $p$. There appears an issue unless $\mathbb{G}_{0}$ has simple closure. To address this issue, we introduce a gluing technique, which turns out to be applicable to various other settings. We view this gluing technique may be the most novel part of this paper. More detailed discussion is given in Section 2.
If we require no restriction on moduli, all known results have restrictions on the Zariski closure. In [BV12] Bourgain-Varjú proved Conjecture 1.1 when the Zariski closure of $\mathbb{G}_{0}$ is $\text{SL}_{d}$, using an extra ingredient from homogeneous dynamics [BFLM11]. Later de Saxcé and He extended this dynamical tool [HdS19], which allowed them to prove Conjecture 1.1 for all $G$ with simple closure [HdS21].
The full generality of Conjecture 1.1, despite abundant evidence, has still remained a technical challenge. In the paper [Gol20], for the purpose of proving Conjecture 1.1 in general, Salehi-Golsefidy conjectured a sum-product phenomenon over finite quotients of rings of algebraic integers. In a separate paper [TZ23], we proved this conjecture:
@@
Suppose $0<\delta\ll 1$, $d$ is a positive integer, and $N_{0}\gg_{d,\delta}1$ is a positive integer. Then there are $0<\varepsilon:=\varepsilon(\delta,d)$ and positive integers $C_{1}=C_{1}(\delta,d),C_{2}=C_{2}(\delta,d),C_{3}=C_{3}(\delta,d)$ such that for any number field $K$ of degree at most $d$ the following holds: Let $\mathcal{O}$ be the ring of integers of $K$. Suppose $\mathfrak{a}$ is an ideal of $\mathcal{O}$ such that $N(\mathfrak{a}):=|\mathcal{O}/\mathfrak{a}|\geq N_{0}$, and suppose $A\subseteq\mathcal{O}$ such that
$|\pi_{\mathfrak{a}}(A)|\geq|\pi_{\mathfrak{a}}(\mathcal{O})|^{\delta}.$
@@
Here, $\pi_{\mathfrak{a}}\text{ (resp. }\pi_{\mathfrak{a}^{\prime}})$ is the reduction map $\mathcal{O}\rightarrow\mathcal{O}/\mathfrak{a}\text{ (resp. }\mathcal{O}\rightarrow\mathcal{O}/\mathfrak{a}^{\prime})$, the set $A^{C_{2}}$ is the $C_{2}$-fold product of the set $A$, and the set $\sum_{C_{3}}A^{C_{2}}$ is the $C_{3}$-fold sum of the set $A^{C_{2}}$.
Roughly speaking, Conjecture 1.1 is a quantitative statement that given a set $A\subset\mathcal{O}$ and an ideal $\mathfrak{a}\subset\mathcal{O}$, if $|\pi_{\mathfrak{a}}(A)|$ is not too small, then modulo an ideal $\mathfrak{a}^{\prime}$ comparable to $\mathfrak{a}$, a sum-product set of $A$ contains a thick arithmetic progression. With this extra ingredient, we can give a new proof of Conjecture 1.1 in the case that $\text{Zcl}(G)=\text{SL}_{2}$, a result due to Bourgain-Varjú, following the spirit of [BG09]. Combining Theorem B with a gluing technique which we think is our main technical contribution, we can prove Conjecture 1.1 for two representative open cases of Conjecture 1.1, which is the main theorem of this paper:
@@
Let $S$ be a finite symmetric set that generates a group $G$ which is a Zariski dense subgroup of ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ or ${\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, then $G$ has super approximation property with respect to all positive integers.
Theorem 1.2 confirms Conjecture 1.1 for the first semisimple but non-simple clousre case, and the first non-semisimple closure case.
@@
Let $S$ be a finite symmetric set in ${\text{SL}}_{2}(\frac{1}{q_{0}}\mathbb{Z})\times{\text{SL}}_{2}(\frac{1}{q_{0}}\mathbb{Z})$ $({\text{SL}}_{2}(\frac{1}{q_{0}}\mathbb{Z})\ltimes(\frac{1}{q_{0}}\mathbb{Z})^{2}$, respectively)))), and $S$ generates a group $G$ which is a Zariski dense in ${\text{SL}}_{2}\times{\text{SL}}_{2},({\text{ASL}}_{2}$, respectively)))), then $G$ has super approximation with respect to $\{q\in\mathbb{Z}_{+},(q,q_{0})=1\}$.
@@ Acknowledgements
We thank Zeev Rudnick, Nicolas de Saxcé, He Weikun and Tran Chieu-Minh for many helpful corrections/suggestions on previous versions of this paper.
@@ 2. Notations and methodology
The unit of any multiplicatively written group is denoted by 1. Occasionally, if a ring structure is present, we denote the additive unit by 0. For given two subsets $A$ and $B$, we denote their product set by $A\cdot B=\{ab|a\in A,b\in B\}$, and their sum set by $A+B=\{a+b|a\in A,b\in B\}$. The $k$-fold product of $A$ is denoted by $A^{k}$, and the $k$-fold sum of $A$ is denoted by $\sum_{k}A$.
If $f$ and $g$ are two complex valued functions on a discrete group $G$, we denote by $f*g$ their convolution
@@
For a prime $p$, we write $p^{n}||q$ if $p^{n}|q$ but $p^{n+1}\nmid q$, and we define $v_{p}(q)=n$. Similarly, for a matrix $M$, $v_{p}(M)=n$ if $p^{n}|M$ but $p^{n+1}\nmid M$. For a prime ideal $\mathcal{P}$ in a ring of integers $\mathcal{O}$, we can define $v_{\mathcal{P}}$ in an analogous way. For two integers $q_{1}$ and $q_{2}$, we write $q_{1}||q_{2}$ if for every $p^{n}||q_{1}$, we also have $p^{n}||q_{2}$. The exact division ``||"``||"` ` | | " can be extended to ideals in a natural way. For a prime ideal $\mathcal{P}$ and a general ideal $\mathfrak{a}$ of a ring $\mathcal{O}$, we write $\mathcal{P}^{n}||\mathfrak{a}$ if $\mathcal{P}^{n}\supset\mathfrak{a}$ but $\mathcal{P}^{n+1}\not\supset\mathfrak{a}$. Similarly, for two ideals $\mathcal{Q}_{1}\supset\mathcal{Q}_{2}$, we write $\mathcal{Q}_{1}||\mathcal{Q}_{2}$ if for all $\mathcal{P}^{n}||\mathcal{Q}_{1}$, we also have $\mathcal{P}^{n}||\mathcal{Q}_{2}$.
For $q=\prod_{i}p_{i}^{n_{i}}\in\mathbb{Z}_{+}$ and $\alpha$ a real positive number, we let $q^{\{\alpha\}}=\prod_{i}p_{i}^{[n_{i}\alpha]}$, where $[n_{i}\alpha]$ is the integer part of $n_{i}\alpha$.
Let $\pi_{q}:\mathbb{Z}\rightarrow q\mathbb{Z}$ be the residue map, which induces residue maps in various other contexts, and we denote them by $\pi_{q}$ as well.
Let $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ or $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, $\Lambda_{q}=\Lambda(\operatorname{mod}q)$ and $\Lambda(q)$ be the kernel of the reduction map $\pi_{q}:\Lambda\rightarrow\Lambda_{q}$. Sometimes we need to reduce the two factors of $\Lambda$ by two different moduli $q_{1},q_{2}$, and we denote this reduction by $\pi_{q_{1},q_{2}}$. Throughout our paper, we fix $\Gamma=\text{SL}_{2}(\mathbb{Z})$.
We use the standard asymptotic notations $O,\Omega,\Theta$ to describe growth of functions. For convenience we also adopt Bourgain’s notations: we write $f(q)<q^{c+}$ to mean $f(q)<q^{c+\epsilon}$ for arbitrarily small $\epsilon$ when $q$ large. Similarly, $f(q)>q^{c-}$ to mean $f(q)>q^{c-\epsilon}$ for arbitrarily small $\epsilon$ when $q$ large.
@@
Then $T_{q}$ is a self adjoint operator on $L^{2}(G_{q})$ with an invariant subspace $l_{0}^{2}(G_{q})$ consisting of functions with average 0. Denote the set of eigenvalues of $T_{q}$ on $l_{0}^{2}(G_{q})$ by $E_{q}$. It follows from Alon and Milman [AM85] that $G$ has super approximation with respect to $\mathbb{Z}_{+}$ if and only if there is some constant $c<1$ independent of $q$ such that $\lambda<c$ for all eigenvalues $\lambda\in E_{q}$. Then following the argument in [BV12] (Proof of Theorem 1, Page 156-158), Theorem 1.2 can be derived from the following proposition:
@@
Let $S$ be a symmetric set in $\Lambda={\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ or ${\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, and assumle that it generates a group $G$ which is Zariski-dense in ${\text{SL}_{2}\times\text{SL}_{2}}$ or ${\text{ASL}_{2}}$. Then for any $\epsilon>0$ there are $10<\delta<$ 1, $C>0$ depending on $\epsilon$ such that the following holds. If $q\in\mathbb{Z}_{+}$ sufficiently large, $A\subset\Lambda$ symmetric, and some integer $l$ satisfying
$\chi_{S}^{(l)}(A)>q^{-\delta},\hskip 14.22636ptl>\delta^{-1}\log q\hskip 14.22636pt\text{ and }\hskip 14.22636pt|\pi_{q}(A)|<|\Lambda_{q}|^{1-\epsilon},$ (2.3)
@@
It is noted that the reduction in [BV12] uses bounds for multiplicities of nontrivial irreducible representations appearing in the regular representation of $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})$. Similar bounds exist for $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})$ and $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\ltimes(\mathbb{Z}/q\mathbb{Z})^{2}$ as well.
In the rest of the paper we will focus on proving Proposition 2.2. We give a sketch here.
@@
We assume (2.4) does not hold for a sufficiently small $\delta$. We will then show a bounded product of $A$ has size at least $|\Lambda_{q}|^{1-\epsilon/2}$, which will force (2.4) to hold afterall.
@@
We assume $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ and we let $\mathbb{P}_{1},\mathbb{P}_{2}$ be the projections of $\Lambda$ to its first and second components. For simplicity we assume each prime divisor of $q$ has sufficiently large exponent. On the first stage, we follow closely the method in [BG08a] and [BG09] and apply Theorem B to show that there is a not-too-small exact divisor $q^{\prime}$ of $q$ and some constant $C$ such that one simple factor, say $\pi_{q^{\prime}}\circ\mathbb{P}_{1}(A^{C})$ contains a large congruence subgroup of $\Gamma_{q^{\prime}}$ (Proposition 5.1). In this process, compared to [BG09], we apply Theorem A to take a shortcut to deduce some results on non-concentration of random walks at Archimedean place.
However, for the purpose of proving Proposition 2.2, we need to take $q^{\prime}$ a very large divisor of $q$, and this leads to the following question:
@@
Suppose for two exact divisors $q_{1},q_{2}$ of $q$, we have a set $B\subset\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ such that $\mathbb{P}_{1}(B),\mathbb{P}_{2}(B)$ are large in $\Gamma_{q_{1}},\Gamma_{q_{2}}$, can we boundedly generate a large subset of $\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ by $B$?
We need to deal with the case $q_{1},q_{2}$ coprime for the above question when growing $q^{\prime}$ to a very large divisor of $q$ for one simple factor, and we also need to consider the case $q_{1},q_{2}$ non co-prime when passing from one simple factor to two simple factors.

### v4

@@ 1. Introduction
For a set $\mathcal{A}$ of positive integers, we say $G$ has super approximation with respect to $\mathcal{A}$ if there is $\varepsilon>0$ such that $\lambda_{0,q}-\lambda_{1,q}>\varepsilon$, $\forall q\in\mathcal{A}$. The existence of a uniform positive spectral gap only depends on $G$ and is independent of the choice of the finite generating set $S$. If $\mathcal{A}=\mathbb{Z}_{+}$, we simply say $G$ has super approximation.
It is well known that lattices in semisimple Lie groups satisfy the super approximation property, a discovery originally due to Margulis [Mar73] and enhanced later by [BS91] and [Clo03], but the involved techniques (spectral method, Property $T$, automorphic forms, etc.) only works for lattices and can not deal with a general non-lattice discrete group, which is also called a thin group. A breakthrough came from Bourgain and Gamburd, who developed the so-called “Bourgain-Gamburd expansion machine”, which is an analytic-combinatorial tool that allows them to prove super approximation property for any Zariski-dense subgroup of $\text{SL}_{2}(\mathbb{Z})$ with respect to prime moduli [BG08b]. A critical ingredient in the proof is Helfgott’s triple product theorem [Hel08]. Since then, there has been a series of papers extending Bourgain-Gamburd’s Theorem to more general groups with respect to more general moduli [BG08a], [BG09], [BGS10], [GV12], [Var12], [BV12], [BGGT15], [PS16], [Gol19], [HdS21].
@@
Let $G<{\text{SL}}_{n}(\mathbb{Z})$ be finitely generated. Then $G$ has the super approximation property if and only if the identity component $\mathbb{G}_{0}$ of the Zariski closure $\mathbb{G}$ of $G$ is perfect, i.e. $\mathbb{G}_{0},\mathbb{G}_{0}]=\mathbb{G}_{0}$.
@@
(Salehi-Golsefidy)[Gol19] Let $G<{\text{SL}}_{n}(\mathbb{Z})$ be finitely generated, then $G$ has the super approximation property with respect to bounded powers of square free integers coprime to $q_{0}$ if and only if the identity component $\mathbb{G}_{0}$ of the Zariski closure $\mathbb{G}$ of $G$ is perfect, i.e. $\mathbb{G}_{0},\mathbb{G}_{0}]=\mathbb{G}_{0}$.
If we require no restriction on moduli, all known results have restrictions on the Zariski closure. In [BV12] Bourgain-Varjú proved Conjecture 1.1 when the Zariski closure of $\mathbb{G}_{0}$ is $\text{SL}_{d}$, using an extra ingredient from homogeneous dynamics [BFLM11]. Later de Saxcé and He extended this dynamical tool [HdS19], which allowed them to prove Conjecture 1.1 for all $G$ with $\mathbb{Q}$-simple closure [HdS21].
The full generality of Conjecture 1.1, despite abundant evidence, has still remained a technical challenge. In the paper [SG20], for the purpose of proving Conjecture 1.1 in general, Salehi-Golsefidy conjectured a sum-product phenomenon over finite quotients of rings of algebraic integers. In a separate paper [TZ23], we proved this conjecture:
@@
Suppose $0<\alpha<1$, $d$ is a positive integer, and $N_{0}\gg_{d,\alpha}1$ is a positive integer. Then there are $0<\varepsilon:=\varepsilon(\alpha,d)$ and positive integers $C_{1}=C_{1}(\alpha,d),C_{2}=C_{2}(\alpha,d),C_{3}=C_{3}(\alpha,d)$ such that for any number field $K$ of degree at most $d$ the following holds: Let $\mathcal{O}$ be the ring of integers of $K$. Suppose $\mathfrak{a}$ is an ideal of $\mathcal{O}$ such that $N(\mathfrak{a}):=|\mathcal{O}/\mathfrak{a}|\geq N_{0}$, and suppose $A\subseteq\mathcal{O}$ such that
$|\pi_{\mathfrak{a}}(A)|\geq|\pi_{\mathfrak{a}}(\mathcal{O})|^{\alpha}.$
@@
Here, $\pi_{\mathfrak{a}}\text{ (resp. }\pi_{\mathfrak{a}^{\prime}})$ is the reduction map $\mathcal{O}\rightarrow\mathcal{O}/\mathfrak{a}\text{ (resp. }\mathcal{O}\rightarrow\mathcal{O}/\mathfrak{a}^{\prime})$, the set $A^{C_{2}}=\{a_{1}a_{2}\cdots a_{C_{2}}:a_{1},\cdots,a_{C_{2}}\in A\}$ is the $C_{2}$-fold product of the set $A$, and the set $\sum_{C_{3}}A^{C_{2}}=\{b_{1}+b_{2}+\cdots+b_{C_{3}}:b_{1},\cdots,b_{C_{3}}\in A^{C_{2}}\}$ is the $C_{3}$-fold sum of the set $A^{C_{2}}$.
Roughly speaking, Conjecture 1.1 is a quantitative statement that given a set $A\subset\mathcal{O}$ and an ideal $\mathfrak{a}\subset\mathcal{O}$, if $|\pi_{\mathfrak{a}}(A)|$ is not too small, then modulo an ideal $\mathfrak{a}^{\prime}$ comparable to $\mathfrak{a}$, a sum-product set of $A$ contains a thick arithmetic progression. With this extra ingredient, we can give a new proof of Conjecture 1.1 in the case that $\text{Zcl}(G)=\text{SL}_{2}$, a result due to Bourgain-Varjú, following the strategy of [BG09]. Combining Theorem B with a gluing technique (Proposition 6.1) which we think is our main technical contribution, we can prove Conjecture 1.1 for two representative open cases of Conjecture 1.1, which is the main theorem of this paper:
@@
Let $S$ be a finite symmetric set that generates a group $G$ which is a Zariski dense subgroup of ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ or ${\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, then $G$ has super approximation with respect to all positive integers.
Theorem 1.2 confirms Conjecture 1.1 for the first semisimple but non-simple closure case, and the first non-semisimple closure case.
@@
Let $S$ be a finite symmetric set in ${\text{SL}}_{2}(\mathbb{Z}[1/q_{0}])\times{\text{SL}}_{2}(\mathbb{Z}[1/q_{0}])$ $({\text{SL}}_{2}(\mathbb{Z}[1/q_{0}])\ltimes(\mathbb{Z}[1/q_{0}])^{2}$, respectively)), and $S$ generates a group $G$ which is a Zariski dense in ${\text{SL}}_{2}\times{\text{SL}}_{2},({\text{ASL}}_{2}$, respectively)), then $G$ has super approximation with respect to $\{q\in\mathbb{Z}_{+},(q,q_{0})=1\}$.
@@ Acknowledgements
We thank Zeev Rudnick, Nicolas de Saxcé, He Weikun, Tran Chieu-Minh and the anonymous referee for many helpful corrections/suggestions on previous versions of this paper.
@@ 2. Notations and methodology
The unit of any multiplicatively written group is denoted by 1. Occasionally, if a ring structure is present, we denote the additive unit by 0. For given two subsets $A$ and $B$, we denote their product set by $A\cdot B=\{ab|a\in A,b\in B\}$, and their sum set by $A+B=\{a+b|a\in A,b\in B\}$. We let $A^{k}=\{a_{1}a_{2}\cdots a_{k}:a_{1},a_{2},\cdots,a_{k}\in A\}$, and $\sum_{k}A=\{a_{1}+a_{2}+\cdots+a_{k}:a_{1},a_{2},\cdots,a_{k}\in A\}$.
If $f$ and $g$ are two complex valued functions on a discrete group $G$, we denote by fgf*g their convolution
@@
For a prime $p$ and an integer $q$, we define $v_{p}(q)=n$ if $p^{n}|q$ but $p^{n+1}\nmid q$. Similarly, for a prime ideal $\mathcal{P}$ and a general ideal $\mathfrak{a}$ of the ring of integers $\mathcal{O}$ of a number field $K$, we define $v_{\mathcal{P}}(\mathfrak{a})=n$ if $\mathcal{P}^{n}\supset\mathfrak{a}$ but $\mathcal{P}^{n+1}\not\supset\mathfrak{a}$. For notational convenience by writing $v_{\mathcal{P}}(a)$ for $a\in\mathcal{\mathcal{O}}$ we mean $v_{\mathcal{P}}((a))$.
The valuation function is naturally extended to matrices: for a matrix $M\in\text{Mat}_{2}(K)$, and a prime ideal $\mathcal{P}$ in $\mathcal{O}$, $v_{\mathcal{P}}(M)=n$ if the minimum of $v_{\mathcal{P}}$-valuations of all coefficients of $M$ is $n$.
The notion “exact division” is extremely helpful for the presentation in this paper. Given a natural number $q$, for each prime $p|q$, we write $p^{n}\|q$ if $v_{p}(q)=n$. Similarly, for two integers $q_{1}$ and $q_{2}$, we write $q_{1}\|q_{2}$ if for every $p|q_{1}$, we have $v_{p}(q_{1})=v_{p}(q_{2})$, and we call $q_{1}$ an exact divisor of $q_{2}$. For two ideals $\mathcal{Q}_{1},\mathcal{Q}_{2}$ of a ring of algebraic integers $\mathcal{O}$, we write $\mathcal{Q}_{1}\|\mathcal{Q}_{2}$ if for all prime ideals $\mathcal{P}\supset\mathcal{Q}_{1}$, we have $v_{\mathcal{P}}(\mathcal{Q}_{1})=v_{\mathcal{P}}(Q_{2})$.
For an ideal $\mathcal{Q}\subset\mathcal{O}$, we say $\mathfrak{q}\in\mathcal{O}$ is a uniformizer of $\mathcal{Q}$ if $\mathfrak{q}\mathcal{O}/\mathcal{Q}^{2}=\mathcal{Q}/\mathcal{Q}^{2}$.
For $q=\prod_{i\in I}p_{i}^{n_{i}}\in\mathbb{Z}_{+}$ and $\alpha$ a positive real number, we let $q^{\{\alpha\}}=\prod_{i}p_{i}^{[n_{i}\alpha]}$, where $n_{i}\alpha]$ is the integer part of $n_{i}\alpha$. For any non-empty subset $U\subset I$, we write $q^{U}=\prod_{i\in U}p_{i}^{n_{i}}$, which is an exact divisor of $q$.
Let $\pi_{q}:\mathbb{Z}\rightarrow\mathbb{Z}/q\mathbb{Z}$ be the residue map, which induces residue maps in various other contexts, and we denote them by $\pi_{q}$ as well.
Let $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ or $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, $\Lambda_{q}=\Lambda(\operatorname{mod}q)$ and $\Lambda(q)$ be the kernel of the residue map $\pi_{q}:\Lambda\rightarrow\Lambda_{q}$. Throughout this paper, we fix $\Gamma=\text{SL}_{2}(\mathbb{Z})$ and $V=\mathfrak{sl}_{2}(\mathbb{Z})=\{\gamma\in\text{Mat}_{2}(\mathbb{Z}):\text{tr}(\gamma)=0\}$.
Sometimes we need to reduce the two factors of $\Gamma\times\Gamma$ by two different moduli $q_{1},q_{2}$, and we denote this residue map by $\pi_{q_{1},q_{2}}$. We also use $\pi_{q_{1},q_{2}}$ to denote the reduction $\Gamma\ltimes\mathbb{Z}^{2}\rightarrow\Gamma_{q_{1}}\ltimes(\mathbb{Z}/q_{2}\mathbb{Z})^{2}$ for $q_{2}|q_{1}$.
We use the standard asymptotic notations $O,\Omega,\Theta$ to describe growth of functions. For two positive functions $f,g$ and a set of parameters, we write $f=O_{X}(g)$ if there is a positive function $C_{1}(X)$ of the set of parameters such that $f\leq C_{1}(X)g$; we write $f=\Omega_{X}(g)$ if there is a positive function $C_{2}(X)$ such that $f\geq C_{2}(X)g$; we write $f=\Theta_{X}(g)$ if $f=O_{X}(g)$ and $f=\Omega_{X}(g)$.
For convenience we also adopt Bourgain’s notations: we write $f(q)<q^{c+}$ to mean $f(q)<q^{c+\varepsilon}$ for arbitrarily small $\varepsilon$ when $q$ large. Similarly, $f(q)>q^{c-}$ means $f(q)>q^{c-\varepsilon}$ for arbitrarily small $\varepsilon$ when $q$ large.
@@
Then $T_{q}$ is a self adjoint operator on $L^{2}(G_{q})$ with an invariant subspace $l_{0}^{2}(G_{q})$ consisting of functions with average 0. Denote the set of eigenvalues of $T_{q}$ on $l_{0}^{2}(G_{q})$ by $E_{q}$. Then following the argument in [BV12] (Proof of Theorem 1, Page 156-158), Theorem 1.2 can be derived from the following proposition:
@@
Let $S$ be a finite symmetric set in $\Lambda={\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ or ${\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, and assumle that it generates a group $G$ which is Zariski-dense in $\text{SL}_{2}\times\text{SL}_{2}$} or $\text{ASL}_{2}$}. Then for any $\varepsilon>0$ there is $0<\delta<1$ depending on $\varepsilon$ such that the following holds. If $q\in\mathbb{Z}_{+}$ sufficiently large , $A\subset\Lambda$ symmetric, and some integer $l$ satisfy
$\chi_{S}^{(l)}(A)>q^{-\delta},\hskip 14.22636ptl>\delta^{-1}\log q\hskip 14.22636pt\text{ and }\hskip 14.22636pt|\pi_{q}(A)|<|\Lambda_{q}|^{1-\varepsilon},$ (2.3)
@@
We recall how to deduce Theorem 1.3 from Proposition 2.2 using the argument from [BV12], Page 156-158. It follows from Alon and Milman [AM85] that $G$ has super approximation with respect to $\mathbb{Z}_{+}$ if and only if there is some constant $c<1$ independent of $q$ such that $\lambda<c$ for all eigenvalues $\lambda\in E_{q}$. Assume $S$ freely generates a group $G$; if not, applying Tits alternative [Tit72] or Proposition 20, [GV12], which generalizes Tits alternative from semisimple groups to perfect groups, we can find a finite symmetric set $S^{\prime}$ which freely generates a subgroup $G^{\prime}<G$, with $G^{\prime}$ Zariski dense in $\Lambda$. Theorem 1.3 for $S^{\prime}$ then implies Theorem 1.3 for $S$.
Let $\lambda$ be an eigenvalue of $T_{q}$ and let $\mu$ be a corresponding eigenfunction. Consider the right regular representation of $\Lambda_{q}$ on $L^{2}(\Lambda_{q})$. Let $\rho$ be the irreducible representation that contains $\mu$. Assume $\rho$ is faithful, i.e., $\rho$ is not induced from a subrepresentation of the right regular representation of $\Lambda_{q^{\prime}}$ for some $q^{\prime}<q$; otherwise, replace $q$ by $q^{\prime}$. According to Proposition 19, [Gol19], if $\lambda$ is outside a finite set $\Sigma$, then $\rho$ has multiplicity at least $|\Lambda_{q}|^{\delta_{0}}$ for some positive $\delta_{0}$, and thus $\lambda$ has multiplicity at least $|\Lambda_{q}|^{\delta_{0}}$. We can now bound $\lambda^{2l}$ by computing the trace of $T^{2l}$ in the standard basis:
$\lambda^{2l}\leq|\Lambda_{q}|^{-\delta_{0}}\text{Tr}(T^{2l})=|\Lambda_{q}|^{1-\delta_{0}}\|\pi_{q}[\chi_{S}^{(l)}]\|_{2}^{2}.$ (2.5)
The Kesten’s bound implies there exists $l_{0}>c_{0}\log q$ for some small $c_{0}>0$, such that
$\|\pi_{q}[\chi_{S}^{(l_{0})}]\|_{2}<|\Lambda_{q}|^{-\varepsilon_{0}}$
for some $\varepsilon_{0}>0$.
If
$\|\pi_{q}[\chi_{S}^{(l_{0})}]\|_{2}\geq|\Lambda_{q}|^{-\frac{1}{2}+\frac{\delta_{0}}{4}},$
then Proposition 2.2 implies
$\|\pi_{q}[\chi_{S}^{(2l_{0})}]\|_{2}<\|\pi_{q}[\chi_{S}^{(l_{0})}]\|_{2}^{1-\delta}$
for some $\delta>0$, with the help of a non-commutative analog of Balog-Szameredi-Gowers theorem due to Bourgain-Gamburd. See [BV12] for details. Keep iterating the convolution until we obtain some $l<C_{0}\log q$ such that
$\|\pi_{q}[\chi_{S}^{(l)}]\|_{2}<|\Lambda_{q}|^{-\frac{1}{2}+\frac{\delta_{0}}{4}}.$ (2.6)
Applying (2.6) to (2.5) leads to $\lambda<e^{-\frac{\delta_{0}}{C_{0}}}$ as desired.
@@
We explain the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$. The case $\Lambda=\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ follows in a similar fashion. Assuming (2.4) does not hold for a sufficiently small $\delta$, we will show a bounded product of $A$ has size at least $|\Lambda_{q}|^{1-\varepsilon/2}$, which will force (2.4) to hold after all.
@@
Let $\mathbb{P}_{1},\mathbb{P}_{2}$ be the projections of $\Lambda$ to its first and second components. If all prime divisors of $q$ have large exponents, we can follow closely the method in [BG08a] and [BG09] and apply Theorem B to show that there is a not-too-small exact divisor $q_{*}$ of $q$ and some constant $C$ such that for one simple factor, say $\pi_{q_{*}}\circ\mathbb{P}_{1}(A^{C})$ contains a large congruence subgroup of $\Gamma_{q_{*}}$ (see Proposition 5.1). In this process, compared to [BG09], we apply Theorem A to take a shortcut to deduce some results on non-concentration of random walks at Archimedean place. However, for the purpose of proving Proposition 2.2, we need to take $q_{*}$ a very large divisor of $q$. This leads to the following question:
Question 2.7.
@@
Suppose for two exact divisors $q_{1}=q^{I_{1}},q_{2}=q^{I_{2}}$ of $q=\prod_{i\in I}p_{i}^{n_{i}}$, $I_{1},I_{2}\subset I$, we have a symmetric set $B\subset\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ such that $\mathbb{P}_{1}(B),\mathbb{P}_{2}(B)$ are large in $\Gamma_{q_{1}},\Gamma_{q_{2}}$, can we boundedly generate a large subset of $\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ by $B$?
Remark 2.8.